OpenStudy (anonymous):

Multiply: 8x^2 divide x^2 - 9 times x^2 + 6x + 9 divide 16x^3

OpenStudy (anonymous):

\[\frac{8x^2}{x^2-9} \times \frac{x^2+6x+9}{16x^3}\] =\[\frac{8x^2}{(x-3)(x+3)} \times \frac{(x+3)^2}{16x^3}\]=\[\frac{x+3}{2x(x-3)}\]

OpenStudy (anonymous):

ty